# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v1

@@ Abstract
Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices. We introduce MetaLint, a new instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications. Unlike conventional approaches that train models on static, rule-based data, MetaLint employs instruction tuning on synthetic linter-generated data to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining. To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize. Our results show that MetaLint improves generalization to unseen PEP idioms, achieving a 70.37% F-score on idiom detection with the highest recall (70.43%) among all evaluated models. It also achieves 26.73% on localization, competitive for its 4B parameter size and comparable to larger state-of-the-art models like o3-mini, highlighting its potential for future-proof code quality analysis.
@@ 1 Introduction
With the rise of Large Language Model (LLM) agents, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent. Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023). However, many of these efforts are hindered by reliance on task-specific training data, often based on narrow or outdated coding best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024). As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024). They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). Ideally, we would develop LLM systems capable of identifying code quality issues without explicit supervision for the target idioms, especially for hard or rare patterns, and that can adapt to evolving best practices over time.
@@
One approach to this problem is to train the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms. For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function. However, as noted in PEP 506 (D’Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against the usage of this module for security reasons, which is often missed by developers as highlighted by accepted answers on forums like StackOverflow. PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom. As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024). Additionally, the issues can be fixed by replacing instances of “bad” idioms with corresponding “good” idioms that align with best practices.
In this work we extended this observation by training the LLM on a higher-level instruction-following task: given a specification of a best-practice code idiom $I$, the model learns to identify and localize code fragments that violate the idiom and could be rewritten to adhere to or incorporate idiom $I$. We refer to this instruction-following task as meta-linting. The core idea is to teach the LLM the underlying structure of the task, namely identifying and locating idioms based on high-level specifications, rather than training it to memorize or learn a static set of best practices. Since the focus is on improving the model’s instruction-following ability rather than tailoring it to a particular data distribution of best practices, this approach can keep up with evolving best practices.
@@
Our approach, MetaLint, is motivated by the observation that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Since meta linting frames each idiom as a separate task, instruction style training is naturally suited to generalize to novel and unseen idioms. Existing linters for programming languages (e.g., Ruff for Python) implement numerous simple rules, or “easy” idioms, that reflect aspects of code quality such as modern syntax, readability, and security. These linters can be used to generate large-scale synthetic data for supervised instruction fine-tuning and serve as verifiers during preference optimization or reinforcement learning. This setup also supports easy-to-hard generalization: training on automatically generated linter data for easy idioms helps build a strong model that can handle harder, more abstract idioms at test time, which go beyond the reach of static analyzers. Similar strategies have shown success in math reasoning tasks (Sun et al., 2024b). To systematically study this generalization, we construct a challenging benchmark of hard idioms based on popular PEPs that introduce high-level constructs. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them to models trained with MetaLint, focusing on whether LLMs can move beyond memorization, generalize from easy to hard idioms, and adapt to novel patterns across code quality domains.
Our key contributions are as follows:
@@
    We introduce MetaLint, an instruction-following framework that leverages synthetic data for easy-to-hard generalization and addresses the limitations of static training used in prior LLM-based code quality approaches.
@@
    We construct a benchmark of challenging, widely relevant idioms inspired by various PEPs, and evaluate state-of-the-art code and reasoning models on it.
@@
    We analyze whether MetaLint promotes adaptive reasoning over memorization and enables generalization from easy to hard idioms.
@@ 2 Related Work
2.1 Code Quality Analysis with Large Language Models
@@
A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis. Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as linters guided by rule descriptions, relying on prompting or fine-tuning to improve rule enforcement. LiteLLM (Fang et al., 2025) leverages LLMs to reduce both the cost and false positive rate of linting for Verilog. Khare et al. (2023) show that LLMs can outperform traditional static analysis tools for detecting certain security-related CWEs, especially when step-by-step reasoning is used. Vijayvergiya et al. (2024) train LLMs to detect and localize violations of best practices, while Rasheed et al. Rasheed et al. (2024) design a multi-agent review pipeline targeting issues like maintainability, efficiency, and bugs. Other works (Jiang et al., 2025b; Yao et al., 2025) use comparative prefix-tuning and reinforcement learning with static analysis–based rewards to promote the generation of higher-quality functionally correct code. Naik et al. (2024) and Jaoua et al. (2025) integrate LLMs with linters to produce more informative code reviews. Finally, CoUpJava (Jiang et al., 2025a) presents a benchmark of Java version–related code upgrade histories, conceptually similar to the hard PEP idiom benchmark we developed for Python. While such efforts highlight the potential of LLMs in code quality tasks, they often target specific rule sets or best practices, necessitating retraining as those evolve. In contrast, our instruction-following meta-linting framework trains models to interpret high-level code quality specifications and perform static analysis to enforce them, enabling broader generalization beyond fixed rules.
2.2 Instruction Following for Generalization
@@
Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning a single fixed task. Prior work shows that exposing models to diverse task instructions allows them to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, acting as a substitute for supervision (Puri et al., 2022) and enabling models to generalize even with minimal labeled data or via pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity is a key driver of generalization, with varied instructions improving performance more than repeated exposure to the same task format (Charton et al., 2024). This phenomenon holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) also performs task-specific instruction following using synthetic data, demonstrating its effectiveness; however, it relies entirely on LLM-generated data and does not incorporate verifiers. Collectively, these results suggest that instruction tuning functions as task-level meta-learning, equipping models with the ability to rapidly adapt to new tasks via natural language. Inspired by this work we treat modeling of specific code quality idioms as individual tasks and propose the MetaLint training framework that generates synthetic data at scale for each meta-task to facilitate cross-idiom generalization. This way the trained model can keep up with new idioms and evolving best practices.
2.3 Easy-to-Hard Generalization
Recent research demonstrates that easy-to-hard generalization improves model performance across mathematical, algorithmic, and coding tasks, motivating its application to code quality analysis. In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025). Several works implement curriculum strategies to structure this progression (Parashar et al., 2025; Zhao et al., 2025), while others emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024). These principles extend beyond math: Sun et al. (2024a) shows that reward models trained on simple code and math problems improve performance on complex ones, and Zhao et al. (2025) further explores generalization in code tasks through curriculum fine-tuning. Multi-task and length generalization studies (Hu et al., 2025) and differentiable programming setups (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code. Zhang et al. (2024) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats. While our approach does not explicitly follow curriculum learning, we draw inspiration from this line of work by training MetaLint on large-scale synthetic data that captures easily detectable code idioms, many of which are directly covered by rule-based linters. We hypothesize that these simple patterns act as a stepping stone, enabling generalization to more complex and abstract unseen PEP idioms.
@@ 3 Method
Figure 1: MetaLint training framework (1/2): Core components of our framework include using linter tools (like Ruff) for 1) Synthetic Data Generation, 2) Supervised Instruction Fine-Tuning (SFT) with synthetic data and a 3) Verifiable Reward Model (also derived from the Ruff linter).
Figure 2: MetaLint training framework (2/2): We use the linter-based verifiable reward model for the final two preference optimization components: 4) Rejection Sampling Direct Preference Optimization (RS-DPO), and, 5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
We propose the MetaLint framework to support the learning of skills essential for code idiom analysis, such as interpreting idiom descriptions in terms of underlying code constructs and performing line-by-line static analysis to detect idiom violations. We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}={D_{I},E_{I}}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization and encourages adaptive reasoning over the prompt’s specification, since flagging violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach enables the LLM to remain flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
@@ 3.1 Synthetic Data Generation
A key advantage of our meta-task formulation is its support for easy-to-hard generalization. We train LLMs on a set of idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$ to improve their ability to flag idiom violations, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by $\mathcal{L}$). Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand, and detect code idioms from high level descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to an untrained checkpoint.
Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For this purpose, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, semantics, and more. We run Ruff on Python source code files $f\in\mathcal{F}$ from the STACK-V2 (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, the Ruff website hosts rule-specific documentation, including descriptions and examples, which we scrape to construct meta-task prompts $M_{I_{\mathcal{L}}}$ for each idiom. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix A.1.1.

### v2

@@ Abstract
Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices. We introduce MetaLint, an instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications. Unlike conventional approaches that train models on static code quality conventions, MetaLint employs instruction tuning on synthetic linter-generated data with dynamic conventions to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining. To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize. Our results show that MetaLint training improves generalization to unseen idioms. Qwen3-4B attains a 70.37% F-score on a manually curated and challenging PEP idiom detection benchmark, achieving the highest recall (70.43%) among all evaluated models. For localization, it reaches 26.73%, which is a strong outcome for its 4B parameter size and comparable to larger state-of-the-art models such as o3-mini, highlighting its potential for future-proof code quality analysis. Furthermore, MetaLint training enables generalization in idiom detection across model families, model scales, synthetic data from diverse linters, and Java idioms, demonstrating the general applicability of our approach. We plan to release our code and data to enable reproducibility and further work.
@@ 1 Introduction
With the rise of Large Language Models (LLM) of code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent Singhal et al. (2024); Zheng et al. (2024). Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Singhal et al., 2024; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023). Several evaluation studies indicate that LLMs struggle with this task Singhal et al. (2024); Zheng et al. (2024), while attempts to improve them through prompting or training are limited by task-specific, static datasets often grounded in narrow or outdated coding practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b). As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024). They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). Ideally, we would develop LLM systems that can identify code quality issues without explicit supervision for target idioms—especially hard or rare patterns—and adapt to evolving best practices over time.
@@
We approach this problem by training the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms. For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function. However, as noted in PEP 506 (D'Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against using this module for security reasons, which is often missed by developers, as highlighted by accepted answers on forums like StackOverflow. PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom. As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024). Additionally, these issues can be addressed by replacing instances of ``bad'' idioms with corresponding ``good'' idioms that align with best practices. Moreover, for this example and similar abstract idioms, constructing a precise rule-based approach is difficult. Simply flagging any use of random.choice, even in non-security-critical scenarios (e.g., randomization in a game engine), could result in a poor user experience. Vijayvergiya et al. (2024) show that LLMs can capture abstract notions of code quality, such as code idioms where building a linter or rule-based approach is challenging, by incorporating semantic reasoning about code and developer intent.
In this work, we train LLMs to recognize code idioms through a higher-level instruction-following task dubbed ``meta-linting'': given a specification of a best-practice code idiom $I$, the model learns to identify and localize non-idiomatic code fragments. Our pipeline is designed to support easy-to-hard generalization (Sun et al., 2024b). The easy cases involve simple idioms that can already be captured by existing linters, while the hard cases correspond to nuanced patterns such as PEP 506, where constructing precise rule-based checks is infeasible. To enable this, we generate synthetic training data for easy idioms using available linters and leverage it to improve performance on harder cases where linter support is lacking. While prior work such as Zhang et al. (2024c; b) has explored automated refactoring of non-idiomatic Python code, including the use of LLMs with prompting, our focus differs in three ways. First, we target challenging idioms beyond the reach of current linters. Second, we train on easy idioms with the goal of transferring detection ability to harder cases. Finally, we emphasize adaptability, aiming for LLMs that can accommodate evolving best practices provided in-context as instructions and examples, rather than memorizing a static rule sets.
@@
To tackle meta-linting, we introduce MetaLint, a training framework motivated by prior work showing that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Since meta-linting treats each idiom as a distinct task or code quality judgment, instruction fine-tuning (IFT) and preference optimization (PO) naturally extend detection ability to novel idioms. Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing simple idioms, which we use both for supervised IFT and as verifiers during PO to improve performance on harder idioms. To systematically study this generalization, we construct a benchmark of challenging idioms derived from popular PEPs introducing high-level constructs. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy idioms.
Our key contributions are:
@@
    We introduce MetaLint, a training framework that leverages instruction following and synthetic data to enable easy-to-hard generalization while remaining adaptable to evolving best practices.
@@
    We construct a benchmark of challenging, broadly relevant code-quality idioms inspired by PEPs to evaluate the extent of easy-to-hard generalization achieved by MetaLint.
@@
    We benchmark state-of-the-art code and reasoning models on our PEP hard-idiom benchmark and compare them against MetaLint-trained models. Our method achieves the highest detection recall and competitive localization scores, even with smaller 4B models and without test-time compute.
  4. 4.
    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), test-time reasoning settings (with and without CoT), and model scales (3B–8B).
@@ 2 Related Work
Code Quality Analysis with Large Language Models. A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis. Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as rule-guided linters via prompting or fine-tuning. While Blyth et al. (2025) proposes a static analysis-driven prompting framework to improve LLM-generated code, Du et al. (2025) conversely uses LLMs to enhance static analysis tools by reducing false-positives. LintLLM (Fang et al., 2025) and (Shin et al., 2025) leverages LLMs for linting of Verilog and Quantum computing code. Khare et al. (2023) show LLMs outperform traditional static analysis tools for security-related CWEs with step-by-step reasoning. Vijayvergiya et al. (2024) train LLMs for best practice violation detection and localization, while Rasheed et al. (2024) design a multi-agent review pipeline for maintainability, efficiency, and bugs. Other works (Jiang et al., 2025b; Yao et al., 2025) use prefix-tuning and reinforcement learning with static analysis–based rewards for higher-quality, functionally correct code generation. Naik et al. (2024) and Jaoua et al. (2025) integrate LLMs with linters to produce more informative code reviews. RIdiom (Zhang et al., 2024c) introduces a rule-based way to identify and refactor non-idiomatic Python code with AST rewrite rules, while Zhang et al. (2024b) combines LLMs and rule-based detectors but doesn't explore nuanced idioms like PEP 506 or training LLMs to keep up with evolving best practices. Finally, CoUpJava (Jiang et al., 2025a) presents Java version upgrade benchmarks, conceptually similar to our hard PEP idiom benchmark for Python. Although prior work demonstrates the potential of LLMs for code quality tasks, it focuses on fixed rule sets or best practices that require retraining as they evolve. In contrast, we train models to interpret high-level specifications and perform static analysis, enabling broader generalization.
@@
Instruction Following for Generalization. Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning fixed tasks. Prior work shows diverse task instructions allow models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, substituting supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity drives generalization, with varied instructions outperforming repeated exposure to identical formats (Charton et al., 2024). This phenomenon holds across domains, including program synthesis where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness, but relying entirely on LLM-generated data without verifiers. These results suggest instruction tuning acts as task-level meta-learning, enabling models to adapt to new tasks through natural language. Building on this we model specific code quality idioms as individual tasks and generate large-scale synthetic data for each meta-task to support cross-idiom generalization. This allows the trained model to keep pace with new idioms and evolving best practices. We also discuss additional related work on easy-to-hard generalization in Appendix B.
@@ 3 Method
We design the MetaLint framework to teach an LLM to operationalize idiom descriptions provided in context, rather than memorizing specific idioms, thereby enabling adaptation to novel idioms at test time. We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization and encourages adaptive reasoning over the prompt’s specification, since flagging violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach enables the LLM to remain flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
@@ 3.1 Synthetic Data Generation
One of the main goals of our meta-task formulation is enabling easy-to-hard generalization. We train LLMs on a set of ``easy'' idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter). Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand and detect code idioms from in-context descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to the untrained model. Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For Python, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, etc., while for Java, we use the PMD static analyzer, which covers 269 idioms as well as some manually written tree-sitter [^1] https://tree-sitter.github.io/tree-sitter/ queries inspired by 8 Java Enhancement Protocols (JEPs) (Table 8). We run Ruff, PMD, and the JEP tree-sitter queries on Python and Java source code files $f\in\mathcal{F}$ from the STACK (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, to automatically build the meta-task instruction prompts $M_{I_{\mathcal{L}}}$ for each idiom, we scrape rule-specific documentation from the Ruff and PMD websites, including descriptions and examples. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix C.1. For the JEP tree-sitter queries, since they are few in number, we manually write the meta-task prompts.